OpenStudy (anonymous):

What are the solutions of 3x2 = 21x? x = 3, x = 7 x = 3, x = −7 x = 0, x = 7 x = 0, x = −7

OpenStudy (anonymous):

\[3x^2-21x=0\] \[3x(x-7)=0\] \[3x=0,x-7=0\] \[x=0,x=7\]